OpenStudy (anonymous):

How would you simplify this given expression, assuming that no variable equals 0?

OpenStudy (anonymous):

.

OpenStudy (anonymous):

\[\left(\frac{3x^{2n}y^2}{x^ny}\right)^3\]

OpenStudy (anonymous):

you can work from the outside in or from the inside out you pick

OpenStudy (anonymous):

How would you do that? Sorry I'm really bad at this.

OpenStudy (anonymous):

first multiply each exponent by 3 to get rid of the 3 outside the parentheses

OpenStudy (anonymous):

don't forget to raise \(3\) to the power of \(3\) as well

OpenStudy (anonymous):

you should get \[\huge\frac{27x^{6n}y^6}{x^{3n}y^3}\]

OpenStudy (anonymous):

next and last step is to subtract the exponents on the like variables

OpenStudy (anonymous):

\[\huge 27x^{6n-3n}y^{6-3}\] is a start